A certain number is added to 20, and the same number is also added to 21; the product of the two sums is 930. What is the number ?
Answer by ankor@dixie-net.com(12701)

A certain number is added to 20,
That's (x+20)
:
and the same number is also added to 21;
That's (x+21)
:
the product of the two sums is 930.
(x+20)*(x+21) = 390
FOIL
x^2 + 41x + 420 = 930
:
x^2 + 41x + 420 - 930 = 0
:
x^2 + 41x - 510 = 0
Factor
(x+51)(x-10) = 0
Two solutions
x = -51
x = +10
:
You can check and see that both solutions are valid